OpenStudy (anonymous):

Help please!!! Half of a number is increased by two and thirds is equal to the number decreased by a quarter. What is the number?

OpenStudy (anonymous):

1/2x+2/3=x-1/4 Try to solve this and tell me what you get

OpenStudy (anonymous):

Oh I'm sorry! I meant two and two thirds!

OpenStudy (anonymous):

Okay, here's the modified equation: 1/2x+8/3=x-1/4 (2 and 2 thirds is the same as 8/3)

OpenStudy (anonymous):

I got 35/6. Is that correct?

OpenStudy (anonymous):

Check your work, numerator wrong

OpenStudy (anonymous):

I'm confused. I keep getting the same answer?

OpenStudy (anonymous):

Show your work, type it, write it (blue button, 'draw')

OpenStudy (anonymous):

1/2x+8/3 = x - 1/4 -1/2-x = -1/4 - 8/3 -1/2x = -35/12 And I got x as 35/6 so I'm not sure what I did wrong...

OpenStudy (anonymous):

So sorry, I was working the problem with 2/3 following the typo above. You are right :)

OpenStudy (anonymous):

Great job @PhiBeeandJay ! Sorry about the confusion with the other user, but you had it right :) I didn't check before, my bad, but hope I helped :)